n the right triangle ABC, AB = 2, BC = 4 and ED is a line parallel to AB. Find the



angle α = angle BAD which minimizes the distance L, where L = AD + ED

1
Expert's answer
2021-12-09T10:13:25-0500

Let "BD=x."


Right triangle "ABD"



"BD=AB\tan \alpha, AD=\dfrac{AB}{\cos \alpha}"

Given "AB = 2, BC = 4."

Substitute



"x=2\tan \alpha, AD=\dfrac{2}{\cos \alpha}"

Since "ED\parallel AB," then right triangles "ABC" and "ABD" are similar.



"\dfrac{ED}{AB}=\dfrac{CD}{BC}=>ED=\dfrac{AB}{BC}\cdot CD""CD=BC-BD=4-2\tan \alpha"

Substitute



"ED=\dfrac{2}{4}(4-2\tan \alpha)=2-\tan \alpha"

Then



"L=AD+ED=\dfrac{2}{\cos \alpha}+2-\tan \alpha,""0\degree<\alpha<90\degree"

Find the first derivative with respect to "\alpha"



"L'_{\alpha}=-\dfrac{2}{\cos ^2 \alpha}(-\sin \alpha)-\dfrac{1}{\cos ^2 \alpha}""=\dfrac{2\sin \alpha-1}{\cos ^2 \alpha}"

Find the critical number(s)



"L'_{\alpha}=0=>\dfrac{2\sin \alpha-1}{\cos ^2 \alpha}=0""\sin \alpha=\dfrac{1}{2}, 0\degree<\alpha<90\degree"

The critical number is "30\degree."

If "0\degree<\alpha<30\degree, L'_{\alpha}<0, L" decreases.

If "30\degree<\alpha<90\degree, L'_{\alpha}>0, L" increases.

The function "L" has a local minimum at "\alpha=30\degree."

Since the function "L" has the only extremum for "0\degree<\alpha<90\degree," then the function "L" has the absolute minimum for "0\degree<\alpha<90\degree" at "\alpha=30\degree."


"\angle BAD=\alpha=30\degree."